OpenStudy (barrelracing):

What is the equation in point-slope form of the line passing through (−2, 0) and (2, 8)? all I need is the formula. I cant remember the formula for this one.

Vocaloid (vocaloid):

we'll need two formulas here, one for the slope and one for the equation slope = (y2-y1)/(x2-x1) equation: y-y1 = m(x-x1)